9/5 is the solution of the equation 4x – 1 = 8.
 9/5 is the solution of the equation 4x – 1 = 8.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The Given statement isFalse
Consider the equation, 4x–1 = 8
Transpose -1 to RHS then it becomes 1.
4x = 8 + 1
4x = 9
x = 9/4

Thanks